# POWSM: A Phonetic Open Whisper-Style Speech Foundation Model

### v1

@@ Abstract
Recent advances in spoken language processing have led to substantial progress in phonetic tasks such as automatic speech recognition (ASR), phone recognition (PR), grapheme-to-phoneme conversion (G2P), and phoneme-to-grapheme conversion (P2G). Despite their conceptual similarity, these tasks have largely been studied in isolation, each relying on task-specific architectures and datasets. In this paper, we introduce POWSM (Phonetic Open Whisper-style Speech Model), the first unified framework capable of jointly performing multiple phone-related tasks. POWSM enables seamless conversion between audio, text (graphemes), and phones, opening up new possibilities for universal and low-resource speech processing. Our model outperforms or matches specialized PR models of similar size (Wav2Vec2Phoneme and ZIPA) while jointly supporting G2P, P2G, and ASR. Our training data, code [^1] https://github.com/espnet and models [^2] https://huggingface.co/espnet/powsm are released to foster open science.
@@
Chin-Jou Li1, Kalvin Chang2, Shikhar Bharadwaj1, Eunjung Yeo3, Kwanghee Choi3, Jian Zhu4, David Mortensen1, Shinji Watanabe1, 1Carnegie Mellon University, 2University of California, Berkeley, 3University of Texas, Austin, 4University of British Columbia,
@@ 1 Introduction
Empirically, POWSM outperforms previous PR models on both in-domain data and out-of-domain languages, achieves low-resource ASR performance comparable to web-scale multilingual foundation models, and can act as speech-grounded P2G and G2P across more than 70 languages. POWSM offers a new unified paradigm for phone-level modeling, paving the way for inclusive and globally accessible speech technologies that transcend language boundaries and resource disparities.
@@
  • We provide POWSM, a large-scale foundation model that achieves state-of-the-art PR performance, and is capable of performing multiple fundamental phone-related tasks. Our model enables seamless conversion between speech, text (graphemes/orthography), and phones.
@@
  • We fully open-source all our data preparation and evaluation scripts, model checkpoints and code to foster open science.
@@ 2 Related Work
Recent speech foundation models such as Whisper (Radford et al., 2023) and OWSM (Peng et al., 2023, 2024) have driven progress in large-scale multilingual ASR and speech translation, but they do not explicitly address phoneme recognition or articulatory-level supervision. Subsequent work (Yusuyin et al., 2025; Fu et al., 2025) showed that incorporating phoneme-level objectives improves ASR for low-resource and long-tailed settings, while outputting phonemes as an intermediate benefited speech translation (Gállego et al., 2025). POWSM extends this line of work by being the first open foundation model jointly trained on phone recognition and related tasks, integrating multilinguality, phonetic supervision, and multi-task scalability within one framework.
@@
G2P & P2G POWSM is the first model capable of both audio-guided G2P and audio-guided P2G. G2P conversion, sometimes called phonemization in the text-to-speech literature, can be accomplished with pronunciation dictionaries (Rudnicky, 1993), rules (Mortensen et al., 2018), WFSTs (Black and Lenzo, 2001), or seq2seq neural methods to choose between different pronunciations of a word in context (Zhu et al., 2022). Text-based G2P, however, still cannot handle phonetic variation, enforcing a one-to-one mapping between orthography and transcription. In contrast, audio-guided G2P can learn to map the different acoustic realizations of a phoneme across varieties of a language to a phone representation Route et al. (2019). In particular, Mak et al. (2025) observed a performance improvement in using audio-guided G2P versus text-based G2P alone for Cantonese. Gao et al. (2024) similarly showed that joint learning of G2P, phone recognition, and forced alignment outperform a G2P teacher model. Similarly, Sun and Richmond (2024) jointly learned G2P and TTS. Compared to G2P, P2G conversion is less studied, with Lauc (2024) training a seq2seq model on 19 million language-grapheme-phoneme triplets.
@@ 3.1 Data preparation
G2P-generated transcriptions have been manually inspected and cleaned. Following Samir et al. (2025), we remove Interlingua and 10 noisy FLEURS languages. Utterances longer than 300 phones are filtered out. IPA sequences are normalized to Unicode NFD (Canonical Decomposition); English G2P sequences are further refined with rule-based corrections to fix voice-onset time issues (see Appendix § A.1).
@@ 3.2 Multitask data format
Our model is trained on four tasks: PR, ASR, and audio-guided G2P and P2G. Each utterance is used once per task, with task-specific formatting as illustrated in Figure˜1, including a text prompt, language token, task token, and target output. We leave the text prompt blank (token <na>) for PR and ASR, and provide graphemes and phones as prompts for G2P and P2G.
@@ 3.3 Training details
POWSM adopts an attention-based encoder-decoder (AED) architecture, which flexibly models output sequences and allows the integration of additional tasks. Specifically, we follow the OWSM v3.1 architecture Peng et al. (2024), which employs an E-Branchformer encoder and a Transformer decoder, consistent with the general encoder-decoder structure of Whisper Radford et al. (2023). The model is trained from scratch using ESPnet Watanabe et al. (2018) with a hybrid CTC/attention loss Watanabe et al. (2017), where we set the ratio $\alpha_{\text{ctc}}$ to 0.3:
$\mathcal{L}=\alpha_{\text{ctc}}\mathcal{L}_{\text{ctc}}+(1-\alpha_{\text{ctc}})\mathcal{L}_{\text{attention}}.$ (1)
@@ 4 Experimental Setup
For unseen languages, we evaluate on three datasets: DoReCo Paschen et al. (2020), VoxAngeles Chodroff et al. (2024), and Tusom2021 Mortensen et al. (2021). DoReCo is a dataset of 50+ languages (with broad transcriptions) intended for documentation of small or endangered languages; we use a 45-language subset. VoxAngeles (Chodroff et al., 2024) is a postprocessed version of the UCLA Phonetics Lab Archive (Ladefoged et al., 2009) containing 95 languages. Tusom is a low-data Tangkhulic language of India not included in the training data. Tusom2021 consists of narrow phonetic transcriptions (unlike the broad transcriptions from G2P on which POWSM was trained) of individual Tusom words. We removed the tones.
@@
We then evaluated our model on in-domain data from IPAPack++, the dataset seen during training. We followed Zhu et al. (2025) in using LibriSpeech for English, AISHELL for Mandarin, and MLS for European languages, and additionally evaluated on IISc-MILE Tamil A et al. (2022) for Tamil and KSC Khassanov et al. (2021) for Kazakh. For ASR and P2G, we evaluate with FLEURS.
@@ 5 Results
We found that POWSM’s performance on PR and ASR tasks is comparable or superior to competitive baselines.
@@ 5.2 POWSM generalizes well to unseen languages
Table 4 reports PFER on datasets with unseen languages and language variation. Results indicate that POWSM achieves strong performance across these datasets, and handles both dialectal and L2 variation effectively. Notably, our method outperforms ZIPA trained on the same data and even exceeds ZIPA trained with extra pseudo-labeled data, achieving the best results on unseen languages while performing three additional tasks. This shows the effectiveness of our multi-task approach. While POWSM lags behind Wav2Vec2Phoneme on socio-phonetic variations, we attribute this to its self-supervised learning with over 60k hours of speech (from wav2vec 2.0 (Baevski et al., 2020)) prior to the supervised learning stage.
@@ 6.1 Behavior of the speech encoder
Figure 2: Validation CER of encoder-CTC during training. Removing suprasegmentals for CTC accelerates convergence and reduces vocabulary size for encoder.
@@
As shown in Table 5, higher CTC decoding weights improve PR performance on out-of-domain data but degrade it on in-domain data, as expected. This echoes Zhu et al. (2025)’s finding that RNN-T (Graves and Jaitly, 2014), an encoder-only speech-to-text model with an autoregressive text prediction network, hurts generalization to unseen patterns of phones (phonotactics). We hypothesize that the decoder is performing implicit language modeling and “smooths” phonetic variation, as Zhu et al. (2025) described.
@@ 6.2 Inspecting Task and Language Tokens
Language token captures phonotactics
@@ 8 Future Work
Finally, the current AED architecture imposes engineering limitations. Inference is significantly slower than with encoder-only models, and the architecture does not easily support tone modeling, limiting its application to tonal languages.
@@ A.3 FLEURS language selection for ASR
A.4 Multi-tasking at Different Scales

### v2

@@ Abstract
Recent advances in spoken language processing have led to substantial progress in phonetic tasks such as automatic speech recognition (ASR), phone recognition (PR), grapheme-to-phoneme conversion (G2P), and phoneme-to-grapheme conversion (P2G). Despite their conceptual similarity, these tasks have largely been studied in isolation, each relying on task-specific architectures and datasets. In this paper, we introduce POWSM (Phonetic Open Whisper-style Speech Model), the first unified framework capable of jointly performing multiple phone-related tasks. POWSM enables seamless conversion between audio, text (graphemes), and phones, opening up new possibilities for universal and low-resource speech processing. Our model outperforms or matches specialized PR models of similar size (Wav2Vec2Phoneme and ZIPA) while jointly supporting G2P, P2G, and ASR. Our training data, code [^1] https://github.com/espnet and model [^2] https://huggingface.co/espnet/powsm are released to foster open science.
@@
Chin-Jou Li1, Kalvin Chang2, Shikhar Bharadwaj1, Eunjung Yeo3, Kwanghee Choi3, Jian Zhu4, David R. Mortensen1, Shinji Watanabe1, 1Carnegie Mellon University, 2University of California, Berkeley, 3University of Texas, Austin, 4University of British Columbia
@@ 1 Introduction
Empirically, POWSM outperforms previous PR models on both in-domain data and out-of-domain languages and achieves low-resource ASR performance comparable to web-scale multilingual foundation models. Moreover, it performs speech-grounded P2G and G2P across more than 70 languages. POWSM offers a new unified paradigm for phone-level modeling, paving the way for inclusive and globally accessible speech technologies that transcend language boundaries and resource disparities.
@@
  • We release POWSM, a large-scale foundation model that achieves state-of-the-art PR performance, and is capable of performing multiple fundamental phone-related tasks. Our model enables seamless conversion between speech, text (graphemes/orthography), and phones.
@@
  • We fully open-source all our data preparation and evaluation scripts, model checkpoint and code to foster open science.
@@ 2 Related Work
Recent speech foundation models such as Whisper (Radford et al., 2023) and OWSM (Peng et al., 2023, 2024) have driven progress in large-scale multilingual ASR and speech translation, but they do not explicitly address phoneme recognition or articulatory-level supervision. Subsequent work (Yusuyin et al., 2025; Fu et al., 2025) showed that incorporating phoneme-level objectives improves ASR for low-resource and long-tailed settings, while outputting phonemes as an intermediate benefited speech translation (gállego2025speechtotexttranslationphonemeaugmentedcot). POWSM extends this line of work by being the first open foundation model jointly trained on phone recognition and related tasks, integrating multilinguality, phonetic supervision, and multi-task scalability within one framework.
@@
G2P & P2G POWSM is the first model capable of both audio-guided G2P and audio-guided P2G. G2P conversion, sometimes called phonemization in the text-to-speech literature, can be accomplished with pronunciation dictionaries (Rudnicky, 1993), rules (Mortensen et al., 2018), WFSTs (Black and Lenzo, 2001), seq2seq models (Zhu et al., 2022), or LLMs (Qharabagh et al., 2025). Text-based G2P, however, still cannot handle phonetic variation, enforcing a one-to-one mapping between orthography and transcription. In contrast, audio-guided G2P can learn to map the different acoustic realizations of a phoneme across varieties of a language to a phone representation Route et al. (2019). In particular, Mak et al. (2025) observed a performance improvement by using audio-guided G2P versus text-based G2P alone for Cantonese. Gao et al. (2024) similarly showed that joint learning of G2P, phone recognition, and forced alignment outperform a G2P teacher model. Similarly, Sun and Richmond (2024) jointly learned G2P and TTS. Compared to G2P, P2G conversion is less studied but can be performed with a seq2seq model (Lauc, 2024) or a finetuned LLM (Ma et al., 2025).
@@ 3.1 Data preparation
G2P-generated transcriptions have been manually inspected and cleaned. Utterances longer than 300 phones are filtered out. IPA sequences are normalized to Unicode NFD (Canonical Decomposition); English G2P sequences are further refined with rule-based corrections to fix voice-onset time issues (see Appendix § A.1).
@@ 3.2 Multitask data format
[…]
@@ 3.3 Training details
POWSM adopts an attention-based encoder-decoder (AED) architecture, which flexibly models output sequences and allows the integration of additional tasks. Specifically, we follow the OWSM v3.1 architecture Peng et al. (2024), which employs an E-Branchformer encoder and a Transformer decoder, consistent with the general encoder-decoder structure of Whisper Radford et al. (2023). The model is trained from scratch using ESPnet Watanabe et al. (2018) with a hybrid CTC/attention loss as in Equation 1 Watanabe et al. (2017), where we set the ratio $\alpha_{\text{ctc}}$ to 0.3.
$\mathcal{L}=\alpha_{\text{ctc}}\mathcal{L}_{\text{ctc}}+(1-\alpha_{\text{ctc}})\mathcal{L}_{\text{attention}}$ (1)
@@ 4 Experimental Setup
For unseen languages, we evaluate on three datasets: DoReCo Paschen et al. (2020), VoxAngeles Chodroff et al. (2024), and Tusom2021 Mortensen et al. (2021). DoReCo is a dataset of 50+ languages (with broad transcriptions) intended for documentation of small or endangered languages; we use a 45-language subset. [^3] We use the same DoReCo subset as Zhu et al. (2025, 2024), listed in § A.6. They removed languages mostly due to licensing issues, while others were not accessible during dataset creation. VoxAngeles (Chodroff et al., 2024) is a postprocessed version of the UCLA Phonetics Lab Archive (Ladefoged et al., 2009) containing 95 languages. Tusom is a low-data Tangkhulic language of India not included in the training data. Tusom2021 (Mortensen et al., 2021) consists of narrow phonetic transcriptions (unlike the broad transcriptions from G2P on which POWSM was trained) of individual Tusom words. We removed the tones.
@@
We then evaluated our model on in-domain data from IPAPack++, the dataset seen during training. We followed Zhu et al. (2025) in using LibriSpeech for English, AISHELL for Mandarin, and MLS for European languages, and additionally evaluated on IISc-MILE Tamil A et al. (2022) for Tamil and KSC Khassanov et al. (2021) for Kazakh.
For ASR and P2G, we evaluate with FLEURS.
@@ 5 Results
We find that POWSM achieves comparable or better performance on PR and ASR than competitive baselines, particularly in low-resource settings; audio-G2P and P2G are discussed in § 6.2. These results suggest that including PR as a pre-training task improves representation generality and reduces the data required to map acoustics to text tokens. In contrast, we find no clear evidence that ASR benefits PR under the current setup (§ A.4, § A.5).
@@ 5.2 POWSM generalizes well to unseen languages
Table 4 reports PFER on datasets with unseen languages and language variation. Results indicate that POWSM achieves strong performance across these datasets, and handles both dialectal and L2 variation effectively. Notably, our method outperforms ZIPA trained on the same data and even exceeds ZIPA trained with extra pseudo-labeled data, achieving the best results on unseen languages while performing three additional tasks. While POWSM lags behind Wav2Vec2Phoneme on socio-phonetic variations, we attribute this to its self-supervised learning with over 60k hours of speech (from wav2vec 2.0 (Baevski et al., 2020)) prior to the supervised learning stage.
@@ 6.1 Behavior of the speech encoder
Figure 2: Validation PER of encoder-CTC during training. Removing suprasegmentals from the training target of encoder accelerates convergence.
@@
As shown in Table 5, higher CTC decoding weights improve PR performance on out-of-domain data but degrade it on in-domain data, as expected. This echoes Zhu et al. (2025)’s finding that RNN-T (Graves and Jaitly, 2014), an encoder-only speech-to-text model with an autoregressive text prediction network, hurts generalization to unseen patterns of phones. We hypothesize that the decoder is performing implicit phonotactic language modeling and “smooths” phonetic variation towards more probabilistically likely phone patterns, as Zhu et al. (2025) described.
@@ 6.2 Inspecting Task and Language Tokens
The language token captures phonotactics
@@ 8 Future Work
Finally, the current AED architecture, although effective for multitasking, imposes certain engineering limitations. Inference is significantly slower than with encoder-only models, and the architecture does not easily support tone modeling, limiting its application to tonal languages.
@@
Acknowledgments
This work used Bridges2 in the PSC and Delta NCSA computing systems through allocation CIS210027 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants. We appreciate the contributions of Kevin Glocker, Shih-heng Wang, Farhan Samir, and Aaricia Herygers during earlier iterations of this work. We are also grateful for feedback from David Harwath, Brendon Boldt, Sanjay Subramanian, Rudy Corona, Anya Ji, Seun Eisape, and Kayo Yin.
@@ A.3 FLEURS language selection for ASR
A.4 ASR Performance on In-Domain Data
We run greedy decoding (ctc=0.0, beam=1) on the test sets of in-domain data. Table 11 shows that ASR performance is reasonably competitive with previous models Radford et al. (2023); Peng et al. (2025) of similar size and architecture despite being trained on much fewer data. The performance gap is smaller for languages that constitute a larger portion of the training data, and larger for mid-frequency languages. Errors often involve substitutions of phonetically similar words or phrases (e.g., “mostly rapscallions as fur” to “mostly wrapped skaggins as fer”).
eng deu nld fra ita spa por pol tam kaz cmn
POWSM 11.5 20.2 13.8 23.6 09.4 25.7 31.0 40.4 23.4 21.9
POWSM-fix 12.2 12.7 19.5 13.9 22.2 11.3 25.5 28.5 32.6 14.8 21.7
Whisper 08.3 11.5 18.2 13.6 21.3 09.1 13.8 12.5 26.6
OWSM 04.7 10.3 15.7 14.1 21.8 07.1 11.5 16.0 14.9
OWSM-CTC 05.1 09.5 15.1 07.8 15.5 05.8 10.3 15.1 18.7
Table 11: WER ($\downarrow$) on test sets; cmn reports CER ($\downarrow$). POWSM-fix is discussed in § A.7. Whisper (Whisper-small) and OWSM (OWSM v4 small) are models of similar size and architecture, trained with 10 times more data. tam and kaz are omitted due to normalization issues; Whisper outputs numerals for cmn, inflating its CER.
A.5 Multi-tasking at Different Scales
@@
A.6 DoReCo (Paschen et al., 2020)
DoReCo consists of the following constituent datasets: (Ozerov, 2022; Riesberg, 2022; Cowell, 2022; Cobbinah, 2022; Vanhove, 2022; Seifart, 2022a; Quesada et al., 2022; Reiter, 2022; Däbritz et al., 2022; Kazakevich and Klyachko, 2022; Hellwig, 2022; Harvey, 2022; Hartmann, 2022; Burenhult, 2022; Kim, 2022; Vydrina, 2022; Gusev et al., 2022; Hellwig et al., 2022; Döhler, 2022; O’Shannessy, 2022a; Bartels and Szczepański, 2022; Haude, 2022; Ponsonnet, 2022; Aznar, 2022; Güldemann et al., 2022; Haig et al., 2022; Garcia-Laguia, 2022; Franjieh, 2022; Ring, 2022; Krifka, 2022; Seifart, 2022b; Witzlack-Makarevich et al., 2022; Xu and Bai, 2022; Forker and Schiborr, 2022; Wegener, 2022; Thieberger, 2022; Schiborr, 2022; Avanzi et al., 2022; Teo, 2022; Gippert, 2022; Bogomolova et al., 2022; Mosel, 2022; Wichmann, 2022; Rose, 2022; Griscom, 2022; Skopeteas et al., 2022; Schnell, 2022; O’Shannessy, 2022b; Michaud, 2022; Skopeteas, 2022; Gipper and Ballivián Torrico, 2022).
A.7 Fixing ASR Text Normalization
After submission, we discovered a text normalization issue specific to Librispeech that degraded its ASR performance. Continuing training with the corrected data for 20 GPUs hours on H100s mostly solved the problem (WER=14.2).
Additionally, we trained a model from scratch with updated data, denoted as POWSM-fix, where ASR transcripts are lowercased and punctuation is removed except for apostrophes and hyphens. As shown in Table 11, ASR performance is similar or improved, while Table 13 shows comparable PR performance on out-of-domain data when decoding with ctc=0.3, beam=1, enabling faster inference. Both checkpoints are publicly available on HuggingFace for reproducibility.
DoReCo VoxA. Tusom. Buckeye DRC-SE L2-ARC EpaDB SO762
POWSM 17.06 17.11 21.96 12.63 18.33 11.32 11.86 17.84
POWSM-fix 19.06 18.80 22.73 13.22 18.84 10.94 10.98 16.56
Table 13: PFER ($\downarrow$) comparison on out-of-domain data.